# AutoPrep: Natural Language Question-Aware Data Preparation with a Multi-Agent Framework

### v1

@@ Abstract.
Answering natural language (NL) questions about tables, which is referred to as Tabular Question Answering (TQA), is important because it enables users to extract meaningful insights quickly and efficiently from structured data, bridging the gap between human language and machine-readable formats. Many of these tables originate from web sources or real-world scenarios, necessitating careful data preparation (or data prep for short) to ensure accurate answers. However, unlike traditional data prep, question-aware data prep introduces new requirements, which include tasks such as column augmentation and filtering for given questions, and question-aware value normalization or conversion. Because each of the above tasks is unique, a single model (or agent) may not perform effectively across all scenarios. In this paper, we propose AutoPrep, a large language model (LLM)-based multi-agent framework that leverages the strengths of multiple agents, each specialized in a certain type of data prep, ensuring more accurate and contextually relevant responses. Given an NL question over a table, AutoPrep performs data prep through three key components. Planner: Determines a logical plan, outlining a sequence of high-level operations. Programmer: Translates this logical plan into a physical plan by generating the corresponding low-level code. Executor: Iteratively executes and debugs the generated code to ensure correct outcomes. To support this multi-agent framework, we design a novel Chain-of-Clauses reasoning mechanism for high-level operation suggestion, and a tool-augmented method for low-level code generation. Extensive experiments on real-world TQA datasets demonstrate that AutoPrep can significantly improve the SOTA TQA solutions through question-aware data preparation.
@@
The source code, data, and/or other artifacts have been made available at https://anonymous.4open.science/r/AutoPrep-Ano.
@@ 1. Introduction
Tabular Question Answering (TQA) refers to the tasks of answering natural language (NL) questions based on provided data tables (Jin et al., 2022; Cheng et al., 2022, 2023; Pal et al., 2023). TQA empowers non-technical users such as domain scientists to easily analyze tabular data and thus has a wide range of applications, including table-based fact verification (Chen et al., 2019; Aly et al., 2021) and table-based question answering (Pasupat and Liang, 2015; Nan et al., 2022).
@@
Figure 2. Examples of Question-Aware Data Preparation for TQA tasks.
Because TQA requires NL understanding and reasoning over tables, state-of-the-art solutions (Wei et al., 2022; Ye et al., 2023; Yao et al., 2023; Cheng et al., 2023; Zhang et al., 2024) naturally rely on large language models (LLMs). These solutions either utilize LLMs as black-boxes or leverage them for code generation. However, question-aware data preparation is not explicitly explored in these approaches, which limits the effectiveness of LLM-based methods.
Figure 1 presents the results of an error analysis of LLM-based code generation (using GPT-4) across two TQA tasks: table-based question answering on the WikiTQ dataset (Pasupat and Liang, 2015) and table-based fact verification on the TabFact dataset (Chen et al., 2019). The analysis reveals that 84% and 76% of the errors stem from insufficient data prep at various granularities (i.e., table-level, column-level, and cell-level) in relation to the NL questions. A similar trend is observed when using LLMs as black-boxes, though those results are not displayed here due to space constraints. These findings underscore the necessity for meticulous data prep to ensure accurate answers.
To better understand these cases, let’s illustrate by examples.
@@
Figure 2 shows several examples of question-aware data prep tasks that are needed for TQA.
(a) Question-Aware Column Augmentation. Figure 2(a) illustrates the necessity of augmenting the tables with new columns, as the information that NL questions require is not explicitly available. For example, to answer $Q_{1}$, a new Country column should be created to augment $T_{1}$, e.g., by extracting country information from Cyclist.
@@
(b) Question-Aware Value Normalization: Figure 2(b) highlights the need for normalizing data types or formats in tables (e.g.,ITA” vs. “Italia” in $T_{4}$, and “19-Oct” vs. “9/14” in $T_{6}$). Note that normalizing tables in TQA requires converting data types or formats based on the need of questions, meaning different questions may result in various operations. For example, as $Q_{5}$ asks for summing numbers in $T_{5}$ but the corresponding columns are of string type, we need to remove commas from numbers and convert strings to integers.
(c) Question-Aware Column Filtering. Figure 2(c) highlights the importance of selecting question-relevant columns from tables, as irrelevant columns add complexity and can lead to errors in code generation. For example, the Result and Record columns in table $T_{7}$ have similar formats and ambiguous meanings, which may mislead the LLM. Additionally, although $T_{8}$ contains 12 columns, only two are relevant to $Q_{8}$. Providing all columns to the LLM can thus result in long-context understanding challenges (Li et al., 2024).
Example 1.1 highlights that different types of data prep operations are needed to satisfy the need of different NL questions.
Given that data prep is a complex challenge without a one-size-fits-all solution, a sensible approach is thus to develop specialized methods w.r.t. each type of data prep operations. These specialized methods, such as data normalization, can involve either LLM-powered code generation or retrieving existing code from data prep libraries. An intelligent module can then automatically determine which operations to perform based on the properties of the question and the corresponding table.
AutoPrep: A Multi-Agent LLM Framework. To address the problem, we propose AutoPrep, a multi-agent LLM framework that automatically prepares tables based on the specific needs of NL questions, eliminating the need for users to specify data prep operations. The core idea is to decompose the data prep task in TQA into sub-tasks, leveraging multiple agents specialized in different areas to ensure accurate and contextually relevant responses.
Given an NL question over a table, AutoPrep performs data prep through the following three core components:
  • Planner: This component is responsible for analyzing the natural language question and the structure of the given table. It generates a logical plan that outlines the sequence of high-level operations required to prepare the data for answering the question. These operations may include tasks such as column augmentation, value normalization, or filtering irrelevant data.
@@
  • Programmer: Once the logical plan is in place, the Programmer translates it into a physical plan by generating low-level executable code. This component is responsible for selecting the appropriate programming constructs (e.g., Python functions or APIs) for each operation, customizing the code to match the table’s structure and data formats.
  • Executor: The Executor is tasked with running the generated code and ensuring its correct execution. It performs iterative execution, where each operation is applied sequentially to the table, and any issues or errors are detected and debugged in real-time.
Technical Challenges. Two key technical challenges arise in our proposed AutoPrep framework.
(C1) Planner: High-Level Operation Suggestion. Even for the same table, different NL questions may require not only distinct data prep operations but also varying sequences of those operations. Thus, suggesting a tailored sequence of high-level operations that meet the specific needs of these questions is a significant challenge.
(C2) Programmer: Low-Level Code Generation. A given high-level operation can have multiple executable code alternatives (e.g., Python functions), and the difference in outcomes between the best and worst options can be substantial. Generating optimized code for each operation that is customized for the tables presents a non-trivial task. For instance, consider the string-to-integer conversion in Example 2.4; failing to account for different value formats can lead to execution errors from a simple type conversion.
Key Ideas. To address the challenge (C1), we propose a Chain-of-Clauses (CoC) based reasoning method for the Planner agent. This method analyzes the NL question and outputs a sequence of high-level desired data prep operations, namely an Analysis Sketch. Based on the Analysis Sketch, the Planner assigns high-level operations to specialized agents for low-level implementation (i.e., code generation).
To tackle challenge (C2), we develop a tool-augmented approach that enhances the LLM’s code generation capabilities by utilizing predefined API functions. We establish a search space of API functions for all high-level operations, design an algorithm to identify the most suitable function for a given operation, and optimize it using a trial-and-error mechanism.
@@
(1) We propose AutoPrep, an LLM-based multi-agent framework for automatic data prep in TQA. We formally define the problem and formalizes typical high-level data prep operations (Section 2), and introduce an overview of the AutoPrep framework (Section 3).
(2) We develop effective techniques for our AutoPrep framework. We propose a novel Chain-of-Clauses reasoning mechanism (Section 4) to suggest high-level operations and a tool-augmented method to generate low-level code (Section 5).
(3) We conduct a thorough evaluation on data prep in TQA (Section 6). Extensive experiments show that AutoPrep achieves new SOTA accuracy on well-adopted TQA datasets. We make the code used in our experiments available on an anonymous GitHub repo [^1] https://anonymous.4open.science/r/AutoPrep-Ano.
2. Preliminaries
This section first introduces Tabular Question Answering (TQA) in Section 2.1 and then formalizes the problem of question-aware data prep for TQA in Section 2.2. Then, we analyze the limitation of existing LLM-based methods in Section 2.3.
@@ 2.1. Tabular Question Answering
Note that TQA often requires complex reasoning over tables. For instance, to answer question $Q_{2}$, we first need to calculate the “GDP growth rate” for all countries, then sort the countries by growth rate, and finally identify the country with the highest GDP growth rate, i.e., CHN.
@@ 2.2. Data Prep for TQA
Unlike traditional data prep, for question-aware data prep for TQA, we need to tailor table $T$ to the specific needs of question $Q$.
Consider Example 2.1: careful data prep, such as extracting countries into a new column and then normalizing them as ISO codes, is crucial for generating an accurate answer. Moreover, it is important to note that different questions may require different data prep tasks, even for the same table. For instance, country extraction may not be necessary if the question is asking which cyclist, instead of country, won the most medals.
Data Prep Operations. To meet the new requirements of data prep for TQA, this paper defines high-level data prep operations (or operations for short) to formalize the question-aware data prep tasks, such as column augmentation, question-aware value normalization or conversion, and column filtering. Formally, an operation, denoted as $o$, encapsulates a specific question-aware data prep task that transforms table $T$ into another table $T^{\prime}$, i.e., $T^{\prime}=o(T)$. Specifically, this paper considers the following three types of operations.
@@
Given a high-level operation $o_{i}$, we define $f_{i}$ as its low-level implementation, either by calling a well-established algorithm from a known Python library or using a customized Python program to meet the requirements of $o_{i}$. For example, to implement the Augment operation $o_{1}$ that augments a Country column in $T_{1}$, a possible implementation $f_{i}$ is to write a customized program by using a specific regular expression as shown in Figure 2(a).
@@
Note that the above three types of operations address most of the errors caused by insufficient data prep, as shown in our error analysis in Figure 1. In fact, the set of data prep operations is extensible, in order to support new or domain-specific data prep tasks.
Question-Aware Data Prep for TQA. Given an NL question $Q$ posed over table $T$, the problem of question-aware data prep for TQA is to generate a sequence of high-level operations $O=\{o_{1},o_{2},\ldots,o_{|O|}\}$ as a logical plan. Then, it generates a physical plan, where each operation $o_{i}$ is implemented by low-level code $f_{i}$, such that these operations transform $T$ into a new table $T^{\prime}$ that meets the needs of $Q$.
@@ 2.3. Limitations of Existing LLM-based Solutions
Naturally, LLMs can be utilized to support question-aware data prep for TQA. A straightforward way is to prompt an LLM with a task description and several demonstrations (i.e., examples), and requires the LLM to generate data prep code, as shown in Figure 3.
However, as shown in Figure 2, the diversity of NL questions necessitates distinct data prep operations. Next, let’s see how a single LLM is used to solve this problem, by an example.
Example 0.
Consider the example table in Figure 3 with columns Date, Cyclist, Medal and Age, along with question: “Which country has the most medals in total in February”. When using few-shot prompting with a task description and a few demonstrations, this strategy may face limitations in both stages of data prep in TQA.
First, as the number of demonstrations is limited, LLMs may struggle to decide which data prep operations are needed for the NL questions, leading to both false negatives (FNs) and false positives (FPs). For example, numeralizing Age in code block ② is an FP, as it is irrelevant to the question. Conversely, the LLM fails to normalize the Date column before extracting the month in code block ③, which is an FN.
Second, due to input token limitations and challenges in long-context understanding (Li et al., 2024), a single LLM may struggle to generate fine-grained code for each operation. For instance, normalizing Medal in code block ③ does not handle all edge cases (e.g., ‘2*’), and the country extraction in code block ⑤ fails to correctly process ‘Dav.ITA’, which is formatted differently from other values.
Example 2.4 shows that, because each of the above tasks is unique, a single model may not perform well across all scenarios. Thus, given that data prep is a complex challenge without a one-size-fits-all solution, a more effective approach is to develop specialized methods for each type. A centralized module can then determine which operations to perform and assign them accordingly.
Figure 3. An LLM-based solution that prompts an LLM with a task description and several demonstrations (i.e., exemplars), and requires the LLM to generate data prep code.

### v2

@@ Abstract.
Answering natural language (NL) questions about tables, known as Tabular Question Answering (TQA), is crucial because it allows users to quickly and efficiently extract meaningful insights from structured data, effectively bridging the gap between human language and machine-readable formats. Many of these tables are derived from web sources or real-world scenarios, which require meticulous data preparation (or data prep) to ensure accurate responses. However, preparing such tables for NL questions introduces new requirements that extend beyond traditional data preparation. This question-aware data preparation involves specific tasks such as column augmentation and filtering tailored to particular questions, as well as question-aware value normalization or conversion, highlighting the need for a more nuanced approach in this context. Because each of the above tasks is unique, a single model (or agent) may not perform effectively across all scenarios. In this paper, we propose AutoPrep, a large language model (LLM)-based multi-agent framework that leverages the strengths of multiple agents, each specialized in a certain type of data prep, ensuring more accurate and contextually relevant responses. Given an NL question over a table, AutoPrep performs data prep through three key components. Planner: Determines a logical plan, outlining a sequence of high-level operations. Programmer: Translates this logical plan into a physical plan by generating the corresponding low-level code. Executor: Executes the generated code to process the table. To support this multi-agent framework, we design a novel Chain-of-Clauses reasoning mechanism for high-level operation suggestion, and a tool-augmented method for low-level code generation. Extensive experiments on real-world TQA datasets demonstrate that AutoPrep can significantly improve the state-of-the-art TQA solutions through question-aware data preparation.
@@
The source code, data, and/or other artifacts have been made available at https://github.com/fmh1art/AutoPrep.
@@ 1. Introduction
Tabular Question Answering (TQA) refers to the task of answering natural language (NL) questions based on provided tables (Jin et al., 2022; Cheng et al., 2022, 2023; Pal et al., 2023). TQA empowers non-technical users such as domain scientists to easily analyze tabular data and has a wide range of applications, including table-based fact verification (Chen et al., 2019; Aly et al., 2021) and table-based question answering (Pasupat and Liang, 2015; Nan et al., 2022). As TQA requires NL understanding and reasoning over tables, state-of-the-art solutions (Wei et al., 2022; Ye et al., 2023; Yao et al., 2023; Cheng et al., 2023; Zhang et al., 2024; Wang et al., 2024b) rely on large language models (LLMs), either utilizing LLMs as black-boxes or leveraging them for code generation.
@@
Figure 2. Examples of Question-Aware Data Preparation.
Question-Aware Data Preparation. As many tables in TQA originate from web sources or real-world data, they demand meticulous data preparation (or data prep) to produce accurate answers. Figure 1 shows an error analysis of an LLM-based approach (using GPT-4) across two TQA tasks: table-based question answering on the WikiTQ dataset (Pasupat and Liang, 2015) and table-based fact verification on the TabFact dataset (Chen et al., 2019). The results show that 84% and 76% of the errors arise from insufficiently addressing the data problems at various granularities (table-level, column-level, and cell-level) in relation to the NL questions. This highlights the critical role of thorough data prep in ensuring accurate responses for TQA.
Figure 2 shows several typical data problems critical to the quality of TQA, which in turn motivate the need of question-aware data preparation.
@@
(a) Column-Level: Missing Column. The example in Figure 2(a) shows the Country column that the NL question $Q_{1}$ requires is not explicitly available.
In this case, an appropriate data prep could address this problem by extracting country information from Cyclist, i.e., augmenting the tables with a new Country column. Clearly, this data augmentation is driven by the specific need of the individual query.
(b) Cell-Level: Inconsistent Data Types or Formats. Figure 2(b) highlights the inconsistency among the data in different columns (e.g.,ITA” vs. “Italia” in $T_{4}$, and “19-Oct” vs. “9/14” in $T_{6}$).
@@
This essentially demands data normalization. This normalization has to be question-aware, because different questions may require different types of normalization. For example, as $Q_{5}$ asks for summing numbers in $T_{5}$ but the corresponding columns are of string type, we need to remove commas from numbers and convert strings to integers.
(c) Table-Level: Irrelevant Columns. As shown in Figure 2(c), some columns in the tables are irrelevant to the questions. For example, $Q_{7}$ does not need the record column in table $T_{7}$. On the other hand, although $T_{8}$ contains 12 columns, only two are relevant to $Q_{8}$.
Filtering out the columns irrelevant to a question is important to TQA, as these irrelevant columns add complexity, potentially misleading LLMs to generate wrong answers. For example, the Result and Record columns in table $T_{7}$ have similar formats and ambiguous meanings, which may confuse the LLM when answering $Q_{7}$. Moreover, this column filtering, avoiding feeding all columns to the LLM, effectively reduces LLM inference cost.
The above examples demonstrate the urgent need of data prep to ensure the quality of TQA. However, blindly applying the existing data prep techniques to address all possible data problems tends to be neither effective nor efficient due to overlooking the specific needs of different NL questions.
In this work, we thus propose to study a new problem, namely question-aware data preparation. Unlike the conventional data prep which prepares tables once at an offline stage to serve many incoming analytics queries (questions), question-aware data prep addresses question specific data problems at online query stage. Once receiving a question, it tailors the tables to the specific need of this question, e.g., performing tasks such as question-specific column augmentation and filtering, as well as question-aware value normalization or conversion.
Key Challenges. Building a system to support question-aware data prep is challenging because of the diversity of both the data problems the tables encounter and the questions. As illustrated in the above examples, different questions may demand different types of data prep, e.g., augmentation, normalization, or filtering. Furthermore, even if facing the same type of data prep problem, different questions may require different ways to handle it. For example, the method that normalizes the date to a unified format is clearly different from that of normalizing strings to integers. Existing LLM-based TQA solutions do not explicitly consider the question-aware data prep problem, thus suffering from the data issues mentioned above. For a detailed analysis, please see Section 2.
AutoPrep: A Hierarchical, Multi-Stage Approach. To address these challenges, we propose AutoPrep, which features two key ideas. First, inspired by the design principles of modern DBMS, in particular the distinction between the logical operations and their physical implementations, AutoPrep separates the high level, logical data prep operations and the concrete data prep methods. That is, it introduces a planning stage to generate a logical plan for each question. The logical plan is composed of a sequence of high-level data prep operations needed by the corresponding questions, including column augmentation, value normalization, or filtering irrelevant data, as shown in Figure 2. AutoPrep then introduces another stage to map the selected logical operations to the corresponding physical operations. In this way, AutoPrep decomposes a data prep task into smaller sub-tasks which are easier to solve than the original complex task. Moreover, it makes AutoPrep extensible, allowing us to easily develop new specialized implementations w.r.t. each type of data prep operations, or supporting new operation types.
Second, unlike conventional DBMS, in AutoPrep the physical operators w.r.t. each logical operator are not developed beforehand. Instead, given a data prep operation in the logical plan, AutoPrep generates a physical implementation specialized to the question on the fly. Taking the question specific need into consideration, this customized implementation thus tends to meet the unique need of different questions.
Driven by the above insight, we design the system into a multi-stage architecture.
The Planning Stage: Unlike DBMS where the logical operators are already available beforehand via SQL queries, AutoPrep has to decide on the appropriate logical data prep operations and in turn form the logical plan by analyzing the NL question and the given table. This is done during the planning stage.
@@
The Programming Stage: This stage translates the logical plan into a physical plan by generating low-level executable code, which is responsible for selecting the appropriate programming constructs (e.g., Python functions or APIs) for each operation, customizing the code to match the table’s structure and data formats, etc.
The Executing Stage: This stage executes the generated code for each operation and returning any errors encountered to the Programming stage for debugging.
We implements this AutoPrep architecture with the popular LLM-based Multi-Agent framework, which uses multiple small, independent agents to collaboratively solve complex problems.
More specifically, we design a Planner agent. Corresponding to the Planning Stage, it suggests a tailored sequence of high-level operations that meet the specific needs of these questions, leveraging the semantics understanding and reasoning ability of LLMs. The key technical idea behind this Planner agent is a novel Chain-of-Clauses (CoC) reasoning method. This method translates the natural language question into an Analysis Sketch, outlining how the table should be transformed to produce the answer and guiding the agent’s reasoning based on this sketch. Compared with the popular Chain of Thoughts (CoT) methods (Wei et al., 2022), which decompose questions into sub-questions, our approach more effectively captures the relationships between questions and tables.
AutoPrep also contains a set of Programmer agent, each of which synthesizes a question specific implementation w.r.t. one logical data prep operation. However, existing LLM-based code synthesis often produces overly generic code that fails to effectively address the heterogeneity challenges of tables. For example, values may have diverse syntactic formats (e.g., “19-Oct” and “9/14” in $T_{6}$) or or semantic representations (e.g., “ITA” and “Italia” in $T_{4}$), making it difficult to generate code tailored to such variations. To address this, we develop a tool-augmented approach that improves the LLM’s code generation capabilities by utilizing predefined API functions. This approach effectively resolves the issue of overly generic code by allowing the LLM to leverage specialized functions tailored to variations in table values. Moreover, corresponding to the Executing stage, we design an Executor agent that executes the generated code to process the table.
@@
(1) We introduce a novel problem of question-aware data preparation for TQA, which is formally defined in Section 2.
(2) We propose AutoPrep, an LLM-based multi-agent framework for question-aware data prep (Section 3). We develop effective techniques in AutoPrep for the Planner agent (Section 4) and the Programmer agents (Section 5).
(3) We conduct a thorough evaluation on data prep in TQA (Section 6). Extensive experiments show that AutoPrep achieves new SOTA accuracy, outperforming existing TQA methods without data prep by 12.22 points on WikiTQ and 13.23 points on TabFact, and surpassing TQA methods with data prep by 3.05 points on WikiTQ and 1.96 points on TabFact.
2. Question-Aware Data Prep for TQA
@@ 2.1. Tabular Question Answering
TQA often requires complex reasoning over tables. For instance, to answer question $Q_{2}$, we first need to calculate the “GDP growth rate” for all countries, then sort the countries by growth rate, and finally identify the country with the highest GDP growth rate, i.e., CHN.
@@ 2.2. Data Prep for TQA
Unlike traditional data prep, for question-aware data prep for TQA, we need to tailor table $T$ to specific needs of question $Q$.
Data Prep Operations. To meet the new requirements of data prep for TQA, this paper defines high-level, logical data prep operations (or operations for short) to formalize the question-aware data prep tasks, such as column augmentation, question-aware value normalization or conversion, and column filtering. Formally, an operation, denoted as $o$, encapsulates a specific question-aware data prep task that transforms table $T$ into another table $T^{\prime}$, i.e., $T^{\prime}=o(T)$. Specifically, this paper considers the following three types of operations.
@@
Given a high-level operation $o_{i}$, we define $f_{i}$ as its low-level implementation, either by calling a well-established algorithm from a known Python library or using a customized Python program to meet the requirements of $o_{i}$.
@@
The above three types of operations address most of the errors caused by insufficient data prep, as shown in our error analysis in Figure 1. In fact, due to the separation of logical operations and physical implementations in AutoPrep, the set of data prep operations can be easily extended to support new data prep tasks.
Question-Aware Data Prep for TQA. Given an NL question $Q$ posed over table $T$, question-aware data prep for TQA is to generate a sequence of high-level operations $O=\{o_{1},o_{2},\ldots,o_{|O|}\}$ as a logical plan. Then, it generates a physical plan, where each operation $o_{i}$ is implemented by low-level code $f_{i}$, such that these operations transform $T$ into a new table $T^{\prime}$ that meets the needs of $Q$.
@@ 2.3. Limitations of Existing LLM-Based Solutions
Figure 3. An LLM-based method with few-shot prompting for question-aware data prep.
A straightforward solution to question-aware data prep is to prompt an LLM to prepare tables, leveraging its ability to interpret the specific requirements of NL questions. Specifically, given the input token constraints of LLMs, it is often infeasible to supply an entire table and request a fully prepared table as output. Therefore, a more effective strategy employs few-shot prompting to generate data prep programs. For instance, consider the table in Figure 3 with columns Date, Cyclist, Medal and Age, and question: “Which country has the most medals in total in February”. The few-shot prompting strategy prompts an LLM with a task description and a few demonstrations, and requests the LLM to generate Python programs as shown in Figure 3.
However, this LLM-based solution may encounter the following limitations when performing question-aware data prep for TQA.
First, given the difficulties of understanding both NL questions and tables, it is challenging to determine which data prep operations are specifically required to meet the needs of each NL question, thus often resulting in false negatives (FNs) and false positives (FPs). Take the results in Figure 3 as an example: converting Age to a numerical format in code block ② is an FP, as it is irrelevant to the question. Conversely, the method fails to normalize the Date column before extracting the month in code block ③, which constitutes an FN, because it does not recognize the inconsistent Date formats.
Second, due to input token limitations and challenges in long-context understanding (Li et al., 2024), it is not easy to fully understand all possible issues in a table, and thus may struggle to generate customized programs to correct issues. For example, in Figure 3, the normalization of Medal in code block ① overlooks certain corner cases (e.g., “2*”), and the country extraction in code block ⑤ fails to handle “Dav.ITA”, which is formatted differently from other values.
Recent methods, such as CoTable (Wang et al., 2024b) and ReAcTable (Zhang et al., 2024), can improve few-shot prompting by employing techniques like Chain-of-Thoughts (CoT) and ReAct. However, these methods remain insufficient to tackle the challenges, as they combine all diverse tasks, such as determining operations and implementing them at various granularities (table-level, column-level, and cell-level), within a single LLM agent. Existing studies (Ishibashi and Nishimura, 2024) have shown that a single LLM agent is often ineffective when tasked with handling a diverse range of operations, due to limited context length in LLMs and decreased inference performance with more input tokens.
Figure 4. A Multi-agent Data Prep Framework for TQA (Left), which could be plugged to any other TQA methods (Right).